# Graph-level Representation Learning with Joint-Embedding Predictive Architectures

### v1

@@ A.1 Experimental setting
Table 5: Performance when using absolute (node-level) vs relative (patch-level) RWSEs.
Patch PE Node-level RWSE Patch-level RWSE
@@
Table 6: Performance when using METIS subgraphs vs. using random subgraphs.
Partitioning method Metis Random
MUTAG 91.25 $\pm$ 5.75 91.58 $\pm$ 5.82
REDDIT-M 56.73 $\pm$ 1.96 56.08 $\pm$ 1.69
IMDB-B 73.68 $\pm$ 3.24 73.52 $\pm$ 3.08
ZINC 0.434 $\pm$ 0.01 0.43 $\pm$ 0.01
Table 7: Results when using the standard self-attention mechanism vs. a graph-specific one.
Attention mechanism Standard Hadamard
MUTAG 91.46 $\pm$ 6.1 91.25 $\pm$ 5.75
REDDIT-M 56.92 $\pm$ 2.09 56.73 $\pm$ 1.96
IMDB-B 73.92 $\pm$ 3.85 73.68 $\pm$ 3.24
ZINC 0.465 $\pm$ 0.01 0.434 $\pm$ 0.01
A.2 Ablation results
Tables 5, 6, and 7 present the results from Figure 3 in tabular format, providing additional clarity.
A.3 Visualizations
In this section, we provide two qualitative examples that give a more visual understanding of how and what Graph-JEPA learns. Firstly, we plot a low-dimensional representation of the high-dimensional latent code obtained from the target encoder when using the Euclidean objective vs using our objective in Figure 4. The modified prediction task greatly alters the geometry and distribution of the latent codes, avoiding the formation of a blob-like structure like in the Euclidean distance task.
Figure 5 depicts how the Graph-JEPA prediction task is framed. In 5a we can see a particular molecule from the ZINC dataset and the corresponding context and target nodes. The target encoder learns a high-dimensional embedding which then can represent the nodes in the 2D unit hyperbola, as shown in 5b. The predictor network tries to guess these coordinates and this training procedure gives rise to the representations learned in Figure 4b.
Figure 4: 3D representation (via t-SNE) of the latent codes used to finetune the linear classifier for the DD dataset. Best viewed in color.
Figure 5: Visualization of a partitioned graph from the ZINC dataset and its corresponding embedding in the 2D unit hyperbola, as detailed in Eq. 4

### v2

@@ 4.1 Experimental setting
4.2 Downstream performance
For the experiments on downstream performance, we follow Suresh et al. (2021) and also report the results of a fully supervised Graph Isomorphism Network (GIN) (Xu et al., 2018), denoted F-GIN in Table 2. We compare Graph-JEPA against four contrastive methods, two generative methods, and one latent self-predictive method (Xie et al., 2022a) (which also regularizes through instance-level reconstruction). As can be seen in Table 2, Graph-JEPA achieves competitive results on all datasets, setting the state-of-the-art as a pretrained backbone on five different datasets and coming second on one (out of eight total). Overall, our proposed framework learns semantic embeddings that work well on different graphs, showing that Graph-JEPA can be utilized as a general pretraining method for graph-level SSL. Notably, Graph-JEPA works well for both classification and regression and performs better than a supervised GIN on all classification datasets. We also provide results with BGRL (Thakoor et al., 2021), a node-level latent self-predictive strategy. We train this model using the official code and hyperparameters and then mean-pool the node representations for the downstream task. The results are underwhelming compared to the models reporting graph-level performance, which is to be expected considering that methods that also perform well on graph-level learning are appropriately designed.
We further explore the performance of our model on the synthetic EXP dataset (Abboud et al., 2020), compared to end-to-end supervised models. This experiment aims to empirically verify if Graph-JEPA can learn highly expressive graph representations (in terms of the commonly used WL hierarchy (Morris et al., 2019)) without relying on supervision. The results in Table 3 show that our model is able to perform much better than commonly used GNNs. Given its local and global exchange of information, this result is to be expected. Most importantly, Graph-JEPA almost matches the flawless performance achieved by He et al. (2023), who train fully supervised.
Table 2: Performance of different graph SSL techniques on various TUD benchmark datasets, ordered by pretraining type: contrastive, generative, and self-predictive. F-GIN is an end-to-end supervised GIN and serves as a reference for the performance values. The results of the competitors are taken as the best values from (Hassani & Khasahmadi, 2020; Suresh et al., 2021; Tan et al., 2023). "-" indicates missing values from the literature. The best results are reported in boldface, and the second best are underlined. For the sake of completeness, we also report the training and evaluation results of GraphMAE on the DD, REDDIT-M5, and ZINC datasets in italics, along with the results of a node-level self-predictive method (BGRL), which does not originally report results on graph-level tasks.
F-GIN 72.39 $\pm$ 2.76 90.41 $\pm$ 4.61 74.87 $\pm$ 3.56 86.79 $\pm$ 2.04 53.28 $\pm$ 3.17 71.83 $\pm$ 1.93 48.46 $\pm$ 2.31 0.254 $\pm$ 0.005
AD-GCL-OPT (Suresh et al., 2021) 73.81 $\pm$ 0.46 89.70 $\pm$ 1.03 75.10 $\pm$ 0.39 85.52 $\pm$ 0.79 54.93 $\pm$ 0.43 72.33 $\pm$ 0.56 49.89 $\pm$ 0.66 $\underline{0.544}$$\pm$ $\underline{0.004}$
GraphMAE (Hou et al., 2022) 75.30 $\pm$ 0.39 88.19 $\pm$ 1.26 74.27 $\pm$ 1.07 88.01 $\pm$ 0.19 46.06 $\pm$ 3.44 $\underline{75.52}$$\pm$ $\underline{0.66}$ $\underline{51.63}$$\pm$ $\underline{0.52}$ 0.935 $\pm$ 0.034
BGRL (Thakoor et al., 2021) 70.99 $\pm$ 3.86 74.99 $\pm$ 8.83 71.52 $\pm$ 2.97 50 $\pm$ 0 20 $\pm$ 0.1 0.5 $\pm$ 0 0.33 $\pm$ 0 1.2 $\pm$ 0.011
LaGraph (Xie et al., 2022a) 75.2 $\pm$ 0.4 90.2 $\pm$ 1.1 78.1 $\pm$ 0.4 90.4 $\pm$ 0.8 56.4 $\pm$ 0.4 73.7 $\pm$ 0.9 - -
Graph-JEPA $\underline{75.67}$$\pm$ $\underline{3.78}$ 91.25 $\pm$ 5.75 78.64 $\pm$ 2.35 91.99 $\pm$ 1.59 56.73 $\pm$ 1.96 73.68 $\pm$ 3.24 50.69 $\pm$ 2.91 0.434 $\pm$ 0.014
Table 3: Classification accuracy on the synthetic EXP dataset (Abboud et al., 2020), which contains 600 pairs of non-isomorphic graphs that are indistinguishable by the 1-WL test. Note that the competitor models are all trained with end-to-end supervision. The best result is reported in boldface, and the second best is underlined. Performances for all competitor models are taken from (He et al., 2023).
Model Accuracy $\uparrow$
GCN (Kipf & Welling, 2016a) 51.90 $\pm$ 1.96
GatedGCN (Bresson & Laurent, 2017) 51.73 $\pm$ 1.65
GINE (Xu et al., 2018) 50.69 $\pm$ 1.39
GraphTransformer (Dwivedi & Bresson, 2020) 52.35 $\pm$ 2.32
Graph-MLP-Mixer (He et al., 2023) 100.00 $\pm$ 0.00
Graph-JEPA 98.77 $\pm$ 0.99
4.3 Exploring the Graph-JEPA latent space
As discussed in Section 3.4, the choice of energy function has a big impact on the learned representations. Given the latent prediction task of Graph-JEPA, we expect the latent representations to display hyperbolicity. The predictor network is linearly approximating the behavior of the unit hyperbola such that it best matches the generated target coordinates (Eq. 6). Thus, the network is actually trying to estimate a space that can be considered a particular section of the hyperboloid model (Reynolds, 1993), where hyperbolas appear as geodesics. We are, therefore, evaluating our energy in a restricted part of hyperbolic space. As mentioned before, we find this task to offer great flexibility as it is straightforward to implement and it is computationally efficient compared to the hyperbolic distance used to typically learn hyperbolic embeddings in the Poincaré ball model (Nickel & Kiela, 2017). Table 4 provides empirical evidence for our conjectures regarding the suboptimality of Euclidean or Poincaré embeddings on 4 out of the 8 datasets initially presented in Table 2, where we make sure to choose different graph types for a fair comparison. The results reveal that learning the distance between patches in the 2D unit hyperbola provides a simple way to get the advantages of both embedding types. Hyperbolic embeddings must be learned in lower dimensions due to stability issues (Yu & De Sa, 2021), while Euclidean ones do not properly reflect the dependencies between subgraphs and the hierarchical nature of graph-level concepts. Our results suggest that the hyperbolic (Poincaré) distance is generally a better choice than the Euclidean distance in lower dimensions, but it is computationally unstable and expensive in high dimensions. The proposed approach provides the best overall results. We provide a qualitative example of how the embedding space is altered from our latent prediction objective in Fig. 3.
Figure 3: 3D t-SNE(Van der Maaten & Hinton, 2008) of the latent representations used to train the linear classifier on the DD dataset. The change in the curvature of the embedding using the Graph-JEPA objective (b.) is noticeable. Best viewed in color.
Table 4: Comparison of Graph-JEPA performance for different distance functions. The optimization for Poincaré embeddings in higher dimensions is problematic, as shown by the NaN loss on the IMDB-B dataset. LD stands for Lower Dimension, where we use a smaller embedding size.
IMDB-B 73.68 $\pm$ 3.24 73.76 $\pm$ 3.46 NaN 72.5 $\pm$ 3.97 73.4 $\pm$ 4.07
Table 5: Total training time and model parameters of MVGRL, GraphMAE, and Graph-JEPA for pretraining (single run) based on the optimal configuration for downstream performance. OOM stands for Out-Of-Memory.
Dataset Model Num. parameters Training time
IMDB MVGRL 3674118 $\sim$ 7 min
GraphMAE 2257193 $\sim$ 1.5 min (1min 36sec)
Graph-JEPA 19219460 $<$ 1min (56 sec)
REDDIT-M5 MVGRL 4198406 OOM
GraphMAE 2784555 $\sim$ 46 min
Graph-JEPA 19245060 $\sim$ 18 min
4.4 Additional insights and ablation studies
Model efficiency.
In an attempt to characterize the efficiency of our proposed model, we perform a simple experimental check. In Table 5, we compare the total training time needed for different Graph-SSL strategies to provide a representation that performs optimally on the downstream task. We show results on the datasets with the largest graphs from Table 2: IMDB and REDDIT-M5. While the runtime is hardware-dependent, all experiments are performed on the same machine. Graph-JEPA displays superior efficiency and promising scaling behavior. The presented runtime is naturally dependent on the self-supervised scheme used in each model, so we do not regard it as a definitive descriptor, but rather an indicator of the potential of fully latent self-predictive models.
MLP parametrization.
Table 6 contains the results of parametrizing the whole architecture, other than the initial GNN encoder, through MLPs. This translates to not using the Attention mechanism at all. For this experiment, and also the following ablations, we consider 4 out of the 8 datasets initially presented in Table 2, making sure to choose different graph domains for a fair comparison. Graph-JEPA still manages to perform well, showing the flexibility of our architecture, even though using the complete Transformer encoders leads to better overall performance and less variance in the predictions.
Table 6: Performance when parametrizing the context and target encoders through MLPs vs using the proposed Transformer encoders.
Dataset Transformer Encoders MLP Encoders
MUTAG 91.25 $\pm$ 5.75 90.5 $\pm$ 5.99
REDDIT-M5 56.73 $\pm$ 1.96 56.21 $\pm$ 2.29
IMDB-B 73.68 $\pm$ 3.24 74.26 $\pm$ 3.56
ZINC 0.434 $\pm$ 0.01 0.472 $\pm$ 0.01
Positional embedding.
Following He et al. (2023), it is possible to use the RWSE of the patches as conditioning information. Formally, let $B\in\{0,1\}^{p\times N}$ be the patch assignment matrix, such that $B_{ij}=1$ if $v_{j}\in p_{i}$. We can calculate a coarse patch adjacency matrix $A^{\prime}=BB^{T}\in\mathbb{R}^{p\times p}$, where each $A_{ij}^{\prime}$ contains the node overlap between $p_{i}$ and $p_{j}$. The positional embedding can then be calculated for each patch by simply using the RWSE described in Eq. 1 on $A^{\prime}$. We test Graph-JEPA with these relative positional embeddings and find that they still provide good performance but consistently fall behind the node-level (global) RWSE that we employ in our formulation (Table 7(b)). An issue of these relative patch RWSEs is that the number of shared neighbors can obscure the local peculiarities of each patch, rendering the context given to the predictor more ambiguous.
Random subgraphs.
A natural question that arises in our framework is how to design the spatial partitioning procedure. Using a structured approach like METIS Karypis & Kumar (1998) is intuitive and leads to favorable results. Another option would be to extract random, non-empty subgraphs as context and targets. As seen in Table 7(a), the random patches also provide strong performance, showing that the proposed JEPA architecture is not reliant on the initial input, as is the case for many methods that rely on data augmentation for view generation (Lee et al., 2022). Even though our results show that using a structured way to extract the patches might not be necessary, random sampling can be problematic for larger graphs. Thus, we advocate extracting subgraphs with METIS as it is a safer option in terms of generalizability across different graphs and the inductive bias it provides.
@@ 4.4 Additional insights and ablation studies
Table 7: (a) Performance when using node-level vs patch-level RWSEs. (b) Performance when extracting subgraphs with METIS vs. using random subgraphs.
Dataset METIS Random
MUTAG 91.25 $\pm$ 5.75 91.58 $\pm$ 5.82
REDDIT-M5 56.73 $\pm$ 1.96 56.08 $\pm$ 1.69
IMDB-B 73.68 $\pm$ 3.24 73.52 $\pm$ 3.08
ZINC 0.434 $\pm$ 0.01 0.43 $\pm$ 0.01
Dataset Node RWSE Patch RWSE
@@
5 Conclusion
In this work, we introduce a new Joint Embedding Predictive Architecture (JEPA) (LeCun, 2022), for graph-level Self-Supervised Learning (SSL). An appropriate design of the model, both in terms of data preparation and pretraining objective, reveals that it is possible for a neural network to self-organize the semantic knowledge embedded in a graph, demonstrating competitive performance in different graph data and tasks. Future research directions include extending the proposed method to nodes and edge-level learning, theoretically exploring the expressiveness of Graph-JEPA, and gaining more insights into the optimal geometry of the embedding space for graph SSL.